OpenStudy (anonymous):

a charge of 1uc is given to one plate of a parallel plate capacitor 0.1uF and a charge of 2uc given to the other plate .find the potential difference develop b/w the plate.

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[C[\text F]=\frac{\Delta Q[\text C]}{\Delta V[\text V]}\]

OpenStudy (anonymous):

nt getting

OpenStudy (unklerhaukus):

\[\Delta V=\frac{\Delta Qd}{\varepsilon A}\]

ganeshie8 (ganeshie8):

Qeffective = 2-1 = 1uc so just for visualizing, if u short the capacitor, .5 uC charge would move from second plate to first plate. so Q on plate we can assume as .5 uC we can use Q=CV now .5 uC = .1uF * V V = .5/.1 = 5V